OpenStudy (anonymous):

can u please help me The sum of the measure of the interior angles of a regular polygon is ten times the sum of the measures of the exterior angles. How many sides does this polygon have?

OpenStudy (anonymous):

sum exterior angkles = 360 so sum int. angles = 3600

OpenStudy (anonymous):

ok but how do find the number of sides it has

OpenStudy (anonymous):

formula is (180- n) / 2 ( i think!!) no - ill have to check that out

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

divide 3600

OpenStudy (anonymous):

180 (n-2) = 3600 180n - 360 = 3600 180n = 3960 n = 3960 / 180 = 22

OpenStudy (anonymous):

ok thanks for showing me how

OpenStudy (anonymous):

no probs plz click good answer to close question - i had to google the formula! lol